n the 4- cable set-up shown, find the tension in each cable to maintain equilibrium of the system.

Cable AD is horizontal

Angle ADE=120°

Angle BDE=105°

Angle CED=105°


1
Expert's answer
2021-09-09T10:49:51-0400


So,


"-AD+BD\cos45°+DE\cos60°=0"

"BD\sin45°-DE\sin60°=1320"

"-DE\cos60°+EC\cos15°=0"

"DE\sin60°+EC\sin15°=1760"


"-AD+BD\cdot0.7071+DE\cdot0.5=0"

"BD\cdot0.7071-DE\cdot0.866=1320"

"-DE\cdot0.5+EC\cdot0.9659=0"

"DE\cdot0.866+EC\cdot0.2588=1760\to"



"AD=3724\ (N)"

"BD=4022\ (N)"

"DE=1760\ (N)"

"EC=911\ (N)" . Answer